# DAPS++: Rethinking Diffusion Inverse Problems with Decoupled Posterior Annealing

### v1

@@ Abstract
From a Bayesian perspective, score-based diffusion solves inverse problems through joint inference, embedding the likelihood with the prior to guide the sampling process. However, this formulation fails to explain its practical behavior: the prior offers limited guidance, while reconstruction is largely driven by the measurement-consistency term, leading to an inference process that is effectively decoupled from the diffusion dynamics. To clarify this structure, we reinterpret the role of diffusion in inverse problem solving as an initialization stage within an expectation–maximization (EM)–style framework, where the diffusion stage and the data-driven refinement are fully decoupled. We introduce DAPS++, which allows the likelihood term to guide inference more directly while maintaining numerical stability and providing insight into why unified diffusion trajectories remain effective in practice. By requiring fewer function evaluations (NFEs) and measurement-optimization steps, DAPS++ achieves high computational efficiency and robust reconstruction performance across diverse image restoration tasks.
@@ 1 Introduction
However, this expected trajectory rarely holds in practice. The data-consistency gradient often dominates the update dynamics, causing each intermediate distribution $\pi(\mathbf{x}_{t}|\mathbf{y})$ to act as an isolated approximation of the posterior rather than part of a coherent sequence of marginals. As a result, the path from $\mathbf{x}_{T}$ to $\mathbf{x}_{0}$ becomes a set of loosely connected estimates rather than a consistent posterior evolution. This reveals that a strictly Bayesian interpretation is insufficient for explaining the empirical behavior of diffusion-based inverse problem solvers. Since diffusion updates and data-consistency updates effectively operate independently, an expectation–maximization (EM)–style viewpoint becomes more suitable: the diffusion stage defines a flat prior space that constrains feasible solutions, while the likelihood gradient $\nabla_{\mathbf{x}_{0}}\log p(\mathbf{y}|\mathbf{x}_{0})$ is refined within this space through an Markov chain Monte Carlo (MCMC)-style update. The two stages are thus naturally decoupled, corresponding to prior initialization followed by likelihood-driven correction.
Building on this perspective, we refine the understanding of why diffusion-based inverse problem solvers are effective and introduce DAPS++. Leveraging the principles of Decoupled Annealing Posterior Sampling (DAPS), our method provides a more efficient and interpretable formulation of diffusion-based Bayesian inference using diffusion models that operate directly in pixel space. Empirically, DAPS++ delivers strong performance across a broad set of inverse problems—including both linear and nonlinear image restoration while offering substantially faster sampling than existing approaches. Compared with DAPS and other diffusion-based inverse problem methods, DAPS++ achieves comparable or superior reconstruction quality while reducing sampling time and neural function evaluations (NFEs) by approximately 90% on both the FFHQ validation set and the ImageNet test dataset.
@@ 2.1 Diffusion Models
Score-based diffusion models [song2019generative, song2021scorebased, song2020improved, karras2022elucidating, ho2020denoising] learn the data distribution $p(\mathbf{x}_{0})$ and its gradient, generating samples by progressively denoising Gaussian-corrupted data. A sequence of isotropic Gaussian perturbations $\mathcal{N}(0,\sigma_{t}^{2}\mathbf{I})$ is applied to the data, inducing a family of noise-conditioned distributions $p(\mathbf{x};\sigma_{t})$. The model is trained to estimate the gradient of the log-density—i.e., the score function—across time steps $t\in[0,T]$, or equivalently across decreasing noise levels $\sigma_{t}$ from $\sigma_{T}=\sigma{\max}$ to $\sigma_{0}=0$. When $\sigma_{\max}$ is sufficiently large, the terminal distribution approaches pure Gaussian noise, $\mathcal{N}(0,\sigma_{\max}^{2}\mathbf{I})$.
During sampling, the learned model approximates the true score of $p_{t}(\mathbf{x};\sigma_{t})$ via $s_{\theta}(\mathbf{x},\sigma_{t})\approx\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x};\sigma_{t})$. Starting from pure noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\sigma_{\max}^{2}\mathbf{I})$, the generative process follows the reverse stochastic differential equation (SDE)
@@ 2.2 Inverse Problems with Diffusion Models
To incorporate the diffusion prior into inverse problems, a standard approach is to form a joint SDE by adding the data-consistency term to Eq. 1, giving
@@
In practice, measurement consistency is defined in terms of the clean likelihood $p(\mathbf{y}|\mathbf{x}_{0})$; as a result, the noised likelihood $p(\mathbf{y}|\mathbf{x}_{t};\sigma_{t})$ is typically inaccessible or ill-defined. Several approaches approximate this quantity [xu2024PnP, song2024Resample, Choi_2021_ICCV] to embed measurement guidance directly into the diffusion process. Among these, Diffusion Posterior Sampling (DPS) [chung2022diffusion] employs Tweedie’s formula to estimate $p(\mathbf{y}|\mathbf{x}_{t})\approx p(\mathbf{y}|\mathbf{x}_{0}=\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}])$, achieving strong empirical performance. However, recent studies show that DPS behaves more like a maximum a posteriori (MAP) estimator than a true posterior sampler [xu2025rethinking, Yang2024Guidance]. This has motivated data-consistency frameworks [song2024Resample, li2024decoupled] that incorporate explicit likelihood optimization into diffusion-based solvers. Nevertheless, approaches that minimize $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t};\sigma_{t})$ often exhibit instability due to overfitting. Another line of work uses SVD-based projections and projection-driven sampling [wang2023DDNM, kawar2022DDRM, chung2022improving] to enforce measurement constraints, while recent decoupled frameworks such as DAPS [zhang2025improving] reformulate Eq. 2 into a two-stage structure that preserves the time-marginal distribution.
However, similarly to DPS, such decoupled methods deviate from posterior sampling as defined in Eq. 2, because they implicitly assume that the time-marginal distributions remain consistent with the true posterior given $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$—an assumption that does not hold in practice. In reality, the two stages function as separate generation and optimization procedures rather than components of a unified posterior sampler, limiting both the theoretical validity and computational efficiency of the framework. This motivates reinterpreting diffusion-based inverse problem solvers primarily as methods for optimizing $p(\mathbf{y}|\mathbf{x}_{0})$ within a diffusion-generated solution space, providing a more accurate perspective for diffusion-based inference.
@@ 3 Method
3.1 Breakdown of Prior Guidance
Motivation. We begin by analyzing the interaction between the prior and likelihood gradients to clarify when and why diffusion priors contribute only marginally to reconstruction. This observation motivates the stage-separation design used in the proposed framework.
To quantify the directional relationship between the prior and likelihood terms, we examine the empirical inner product
@@ 3.1 Breakdown of Prior Guidance
where $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t};\sigma_{t})$ and $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t};\sigma_{t})$ denote the likelihood and prior gradients at noise level $\sigma_{t}$, respectively. Under high-noise conditions, the expected value of this inner product is close to zero, implying that the two gradients are effectively uncorrelated. This orthogonality helps explain why methods such as DPS can combine these terms without destructive interference (see Section 3.4).
@@
where $\kappa_{t}$ measures the dominance of the data-consistency gradient relative to the prior gradient. Under the Gaussian noise model $p(\mathbf{y}|\mathbf{x}_{0})=\mathcal{N}(\mathcal{A}(\mathbf{x}_{0}),\gamma^{2}\mathbf{I})$, the likelihood term at noise level $\sigma_{t}$ may be approximated by
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t};\sigma_{t})\simeq-\frac{1}{\gamma^{2}}\,\nabla_{\mathbf{x}_{t}}\left\|\mathbf{y}-\mathcal{A}\big(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})\big)\right\|^{2}$ (5)
@@
Figure 1: (a) Evolution of the gradient ratio $\kappa_{t}$ with respect to noise level and its relative error during a Gaussian-blur iteration, illustrating that the data-consistency gradient dominates throughout the optimization process. (b) Comparison between DAPS results using the prior term and using only the data term across different numbers of annealing steps, showing that the prior contributes minimally and mainly shapes the time-marginal distribution.
Together with the criterion $\kappa_{t}\!\gg\!1$, as shown in Fig. 1(a), the analysis indicates that the diffusion prior provides only a negligible contribution to the optimization, supporting its removal in the DAPS++ framework. A corresponding $L$-Lipschitz argument in the Appendix further shows that the prior gradient $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t};\sigma_{t})$ within the time-marginal formulation makes only a limited impact on the update. In practice, reconstruction is driven almost entirely by the data-consistency term, and as confirmed in Fig. 1(b), omitting the prior component in the time-marginal yields nearly identical reconstruction quality, validating its minimal influence.
3.2 Complete Stage Separation via EM Reformulation
Under this observation, decoupled noise-annealing approaches do not strictly follow the time-marginal distribution, especially when initialized from high-noise conditions. Nevertheless, they perform well empirically, as the diffusion model still produces high-quality reconstructions from the measurements. To explain this behavior, we reinterpret the process as two fully independent stages inspired by an expectation–maximization (EM) viewpoint, yielding a complete separation between diffusion-based initialization and posterior refinement, as illustrated in Fig. 2.
Figure 2: Diagram of the DAPS++ framework. The E-step provides the initial state and constructs the constrained optimization space $p(\mathbf{x}_{0})$ for data-driven MCMC refinement, while the M-step optimizes within this space under measurement guidance. The two steps are fully decoupled, and after each M-step, controlled noise is re-injected by the diffusion process to initiate the next E-step.
Expectation step. Solving the diffusion SDE produces samples consistent with the learned prior $p(\mathbf{x}_{0})$, providing an initialization whose mean and variance are close to the target distribution. Because high-noise initialization randomizes the state far from the data manifold, using higher-order solvers or more diffusion steps in DAPS typically does not further improve pixel-domain initialization. Tweedie’s formula offers a principled estimator of the conditional mean:
@@ 3.2 Complete Stage Separation via EM Reformulation
where $s_{\theta}(\mathbf{x}_{t},\sigma_{t})$ denotes the learned score function. To balance efficiency and fidelity, DAPS++ introduces a noise threshold $\bar{\sigma}$: for $\sigma_{t}>\bar{\sigma}$, Tweedie denoising (Eq. 6) yields fast initialization; for $\sigma_{t}\leq\bar{\sigma}$, a single higher-order ODE step is used to preserve fine details at minimal cost.
Maximization step. This step refines the estimate by performing a likelihood–driven update to enforce measurement consistency within the diffusion–constrained domain. Since the diffusion model initializes samples close to the high–density region of $p(\mathbf{x}_{0})$ (See Appendix about manifold–proximity analysis), the search space is already aligned with the target posterior, allowing Markov chain Monte Carlo (MCMC) to converge with substantially fewer iterations. Following the unadjusted Langevin algorithm (ULA), the update rule is
@@
where $\eta$ is the step size and $\boldsymbol{\epsilon}_{j}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$ denotes Gaussian perturbation for stochastic exploration.
@@
In standard MCMC, the prior gradient term $\nabla_{\mathbf{x}_{0}^{(j)}}\log p(\mathbf{x}_{0}^{(j)})$ serves as a regularization component. However, because diffusion initialization places $\mathbf{x}_{0}^{(j)}$ in a region where the score magnitude satisfies $\|\nabla\log p(\mathbf{x}_{0})\|\!=\!O(\varepsilon)$, this contribution becomes numerically negligible. We therefore omit it and obtain a likelihood–driven refinement step that remains stable and computationally efficient, since only a few informed updates are required within the restricted domain.
Although this refinement drives the estimate toward the clean target, it reduces variability associated with noise–corrupted or unobserved components. To restore this variability, each M–step is followed by a re–noising stage that returns to the E–step, reintroducing structured uncertainty according to
$\mathcal{Q}_{t}=\mathbb{E}_{\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t})}\Big[\log p\big(\mathbf{x}_{0},\,\mathbf{x}_{t}\mid\mathbf{y}\big)\Big],$ (8)
which preserves the statistical characteristics of the diffusion prior while maintaining consistency with the measurement constraints.
Unlike a maximum a posteriori (MAP) approach that seeks a single optimum $\arg\max_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t};\sigma_{t})$, our method performs stochastic sampling within the diffusion-defined latent space, thereby exploring the posterior distribution rather than collapsing to one mode. To further improve convergence stability, we adopt an EDM-inspired annealing schedule in which the effective noise or step size decays polynomially (with exponent in the range [4,7][-4,-7]), enabling smoother and more stable refinement under low-noise conditions.
3.3 DAPS++ for General Inverse Problems
The complete DAPS++ framework consists of two sequential stages that together form a fully decoupled pipeline.
Stage 1: Diffusion-Based Initialization. Starting from pure noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\sigma_{\max}^{2}\mathbf{I})$, we run the reverse diffusion process without any likelihood guidance to obtain $\hat{\mathbf{x}}_{0}$ using Eq. 6 as noise scheduler smaller than $\bar{\sigma}$. This stage serves as a one-time initialization that approximates the manifold of the clean data distribution. The resulting $\hat{\mathbf{x}}_{0}$ provides a sample within a flat prior space that captures the global structure of the signal and is passed directly to Stage 2 without invoking any additional prior information, ensuring complete decoupling between the two stages. Stage 2 only requires $\hat{\mathbf{x}}_{0}\in\mathcal{M}_{p(\mathbf{x}_{0})}$, since the subsequent MCMC refinement is independent of the specific initialization path.
Stage 2: Likelihood-Driven MCMC Refinement. Given the initialization $\hat{\mathbf{x}}_{0}$, we perform independent MCMC sampling driven solely by the likelihood term under the Gaussian measurement model. The update rule is
@@ 3.3 DAPS++ for General Inverse Problems
$\mathbf{x}_{0}^{(j+1)}=\mathbf{x}_{0}^{(j)}-\frac{\eta}{\gamma^{2}}\,\nabla_{\mathbf{x}_{0}^{(j)}}\left\|\,\mathbf{y}-\mathcal{A}\!\left(\mathbf{x}_{0}^{(j)}\right)\right\|^{2}+\sqrt{2\eta}\,\boldsymbol{\epsilon}_{j}$ (9)
where $\gamma$ controls the measurement noise level and $\boldsymbol{\epsilon}_{j}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ adds stochastic exploration. This refinement stage updates the sample strictly according to the measurement likelihood, allowing inference to be guided directly by the observation model while maintaining numerical stability.
While DAPS performs alternating E–M updates at every annealing diffusion timestep to preserve the time-marginal distribution, DAPS++ completes Stage 1 entirely before moving to Stage 2, thereby removing the need to balance prior and likelihood gradients during diffusion sampling. This full separation not only reduces the computational cost associated with repeated diffusion evaluations but also improves interpretability by isolating the generative prior (diffusion) from the data-consistency refinement (MCMC). A detailed algorithmic description and pseudocode are provided in the Appendix.

### v2

@@ Abstract
From a Bayesian perspective, score-based diffusion solves inverse problems through joint inference, embedding the likelihood with the prior to guide the sampling process. However, this formulation fails to explain its practical behavior: the prior offers limited guidance, while reconstruction is largely driven by the measurement-consistency term, leading to an inference process that is effectively decoupled from the diffusion dynamics. We show that the diffusion prior in these solvers functions primarily as a warm initializer that places estimates near the data manifold, while reconstruction is driven almost entirely by measurement consistency. Based on this observation, we introduce DAPS++, which fully decouples diffusion-based initialization from likelihood-driven refinement, allowing the likelihood term to guide inference more directly while maintaining numerical stability and providing insight into why unified diffusion trajectories remain effective in practice. By requiring fewer function evaluations (NFEs) and measurement-optimization steps, DAPS++ achieves high computational efficiency and robust reconstruction performance across diverse image restoration tasks.
@@ 1 Introduction
However, this expected trajectory rarely holds in practice. The gradient of data likelihood gradient often dominates the update dynamics, causing each intermediate distribution $\pi(\mathbf{x}_{t}|\mathbf{y})$ to act as an isolated approximation of the posterior rather than part of a coherent sequence of marginals. As a result, the path from $\mathbf{x}_{T}$ to $\mathbf{x}_{0}$ becomes a set of loosely connected estimates rather than a consistent posterior evolution. This reveals that a strictly Bayesian interpretation is insufficient for explaining the empirical behavior of diffusion-based inverse problem solvers. This reveals that the diffusion prior and the data-consistency update serve fundamentally different roles: the diffusion stage provides a structured initialization that constrains the feasible solution space to lie near the data manifold, while the likelihood gradient refines this initialization to satisfy the measurement. The two stages are thus naturally decoupled, corresponding to prior-driven initialization followed by likelihood-driven correction.
Building on this perspective, we introduce DAPS++. Leveraging the principles of Decoupled Annealing Posterior Sampling (DAPS), our method provides a more efficient and interpretable formulation of diffusion-based Bayesian inference using diffusion models that operate directly in pixel space. Empirically, DAPS++ delivers strong performance across a broad range of inverse problems, including both linear and nonlinear image restoration, while offering substantially faster sampling than existing approaches. Compared with DAPS and other diffusion-based inverse problem methods, DAPS++ achieves comparable or superior reconstruction quality while reducing sampling time and neural function evaluations (NFEs) by approximately 90%on both the FFHQ validation set and the ImageNet test dataset.
@@ 2.1 Diffusion Models
Score-based diffusion models [song2019generative, song2021scorebased, song2020improved, karras2022elucidating, ho2020denoising] learn the data distribution $p(\mathbf{x}_{0})$ and its gradient, generating samples by progressively denoising Gaussian corrupted data. A sequence of isotropic Gaussian perturbations $\mathcal{N}(0,\sigma_{t}^{2}\mathbf{I})$ is applied to the data, inducing a family of noise-conditioned distributions $p(\mathbf{x};\sigma_{t})$. The model is trained to estimate the gradient of the log-density as a score function across time steps $t\in[0,T]$, or equivalently across decreasing noise levels $\sigma_{t}$ from $\sigma_{T}=\sigma{\max}$ to $\sigma_{0}=0$. When $\sigma_{\max}$ is sufficiently large, the terminal distribution approaches pure Gaussian noise, $\mathcal{N}(0,\sigma_{\max}^{2}\mathbf{I})$.
During sampling, the learned model approximates the true score of $p_{t}(\mathbf{x};\sigma_{t})$ via $s_{\theta}(\mathbf{x},\sigma_{t})\approx\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x};\sigma_{t})$. Starting from pure noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\sigma_{\max}^{2}\mathbf{I})$, the generative process follows the reverse stochastic differential equation (SDE) [lu2022dpm, song2019generative]
@@ 2.2 Inverse Problems with Diffusion Models
To incorporate the diffusion prior into inverse problems, a standard approach is to form a joint SDE by adding the data-consistency term to Eq.˜1, giving
@@
In practice, measurement consistency is defined in terms of the clean likelihood $p(\mathbf{y}|\mathbf{x}_{0})$; as a result, the noised likelihood $p(\mathbf{y}|\mathbf{x}_{t};\sigma_{t})$ is typically inaccessible or ill-defined. Several approaches approximate this quantity [xu2024PnP, song2024Resample, Choi_2021_ICCV] to embed measurement guidance directly into the diffusion process. Among these, Diffusion Posterior Sampling (DPS) [chung2022diffusion] employs Tweedie’s formula to estimate $p(\mathbf{y}|\mathbf{x}_{t})\approx p(\mathbf{y}|\mathbf{x}_{0}=\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}])$, achieving strong empirical performance. However, recent studies show that DPS behaves more like a maximum a posteriori (MAP) estimator than a true posterior sampler [xu2025rethinking, Yang2024Guidance]. This has motivated data-consistency frameworks [song2024Resample, li2024decoupled] that incorporate explicit likelihood optimization into diffusion-based solvers. Nevertheless, approaches that minimize $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t};\sigma_{t})$ often exhibit instability due to overfitting. Another line of work uses SVD-based projections and projection-driven sampling [wang2023DDNM, kawar2022DDRM, chung2022improving] to enforce measurement constraints, while recent decoupled frameworks such as DAPS [zhang2025improving] reformulate Eq.˜2 into a two-stage structure that preserves the time-marginal distribution.
However, similar to DPS, such decoupled methods deviate from posterior sampling as defined in Eq.˜2, because they implicitly assume that the time-marginal distributions remain consistent with the true posterior given $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$, which does not hold in practice. In reality, as we show below, the two stages operate as separate generation and optimization procedures rather than components of a unified posterior sampler, limiting both the theoretical validity and computational efficiency of the framework. This observation motivates reinterpreting diffusion-based inverse problem solvers primarily as methods for optimizing $p(\mathbf{y}|\mathbf{x}_{0})$ within a diffusion-generated solution space, offering a more accurate perspective on diffusion-based inference.
@@ 3 Method
3.1 The Diffusion Prior as a Warm Initializer
We begin by analyzing the interaction between the prior and likelihood gradients to clarify the functional role of the diffusion prior in inverse problem solving.
To quantify the directional relationship between the two terms, we examine the empirical inner product
@@ 3.1 The Diffusion Prior as a Warm Initializer
where the two factors denote the likelihood and prior score at noise level $\sigma_{t}$, respectively. Under high-noise conditions, this inner product is empirically close to zero, indicating that the two gradients are effectively orthogonal. This orthogonality helps explain why methods such as DPS can combine both terms without destructive interference (see Sec.˜3.4).
@@
Under the Gaussian noise model $p(\mathbf{y}|\mathbf{x}_{0})=\mathcal{N}(\mathcal{A}(\mathbf{x}_{0}),\gamma^{2}\mathbf{I})$, the likelihood gradient at noise level $\sigma_{t}$ can be approximated as
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t};\sigma_{t})\simeq-\frac{1}{\gamma^{2}}\,\nabla_{\mathbf{x}_{t}}\left\|\mathbf{y}-\mathcal{A}\big(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})\big)\right\|^{2},$ (5)
@@
Figure 1: (a) Evolution of gradient ratio $\kappa_{t}$ vs. noise level during a Gaussian-blur iteration, showing data-consistency gradient dominates throughout. (b) Reconstruction quality of DAPS when driven by the prior (time-marginal) vs. the data likelihood term only across annealing steps, showing that data likelihood alone accounts for nearly all reconstruction quality.
Figure 2: Diagram of the DAPS++ framework. Stage 1: Initialization constructs the constrained optimization space $p(\mathbf{x}_{0})$. Stage 2: Refinement optimizes under measurement guidance. The two stages are fully decoupled, with noise re-injected between steps.
As shown in Fig.˜2(a), $\kappa_{t}\gg 1$ throughout the diffusion process. A Lipschitz-bound analysis in the Appendix (Sec. 6.1) formalizes this: under standard regularity assumptions, $\kappa_{t}\gtrsim\frac{\sigma_{\min}^{+}(\mathcal{A})}{\gamma^{2}C}\sigma_{t}|\mathbf{r}_{t}|$, which grows with the noise-to-measurement ratio $\sigma_{t}/\gamma$. This does not imply that the diffusion prior is unnecessary; instead, it is essential for placing the initial estimate near the data manifold. The analysis shows that once initialization is achieved, the prior gradient contributes negligibly to subsequent updates: it is small in magnitude ($\kappa_{t}\gg 1$) and approximately orthogonal to the likelihood gradient ($A_{t}\approx 0$), meaning it neither accelerates nor redirects the measurement-driven correction. As confirmed in Fig.˜2(b), removing the prior term from the time-marginal update and relying solely on the data-consistency term produces nearly identical reconstruction quality. This demonstrates that the continuous prior guidance used in existing methods can be fully replaced by a one-time diffusion-based initialization.
3.2 Complete Stage Separation: Initialization and Refinement
The observation above suggests that decoupled noise-annealing approaches do not strictly follow the time-marginal distribution, especially when initialized from high-noise conditions. Nevertheless, they produce strong reconstructions empirically because the diffusion model provides a high-quality initialization that the likelihood gradient can then refine. This motivates a complete separation between the two stages, as illustrated in Fig.˜2.
Stage 1: Diffusion Initialization. We start from $\mathbf{x}_{T}\sim\mathcal{N}(0,\sigma_{\max}^{2}\mathbf{I})$ and run the reverse diffusion process without likelihood guidance. A key design question is what form the initialization should take. In the EDM noise schedule, the probability flow ODE is
$d\mathbf{x}=-\dot{\sigma}\sigma\,\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma)\,dt,$ (6)
and a single Euler step from $\mathbf{x}_{t}$ recovers exactly Tweedie’s formula:
@@ 3.2 Complete Stage Separation: Initialization and Refinement
where $s_{\theta}(\mathbf{x}_{t},\sigma_{t})$ denotes the learned score function. As a first-order mean estimator, Tweedie’s formula produces an initialization near the data manifold that is smooth and fast to compute. Higher-order ODE solvers follow the diffusion trajectory more faithfully but do not improve Stage 2 convergence: what matters for refinement is proximity to the manifold, not precision of the sampling path, since the likelihood gradient handles the remaining correction. Appendix Fig. 6 validates this by comparing initialization strategies across different sampling method settings, confirming that a single Tweedie step at an appropriate $\sigma$} provides a more effective starting point than ODE-based alternatives.
To balance efficiency and fidelity, DAPS++ introduces a noise threshold $\bar{\sigma}$: for $\sigma_{t}>\bar{\sigma}$, Tweedie denoising (Eq.˜7) provides fast initialization; for $\sigma_{t}\leq\bar{\sigma}$, a single higher-order ODE step (RK4) preserves fine details at minimal additional cost. The resulting $\hat{\mathbf{x}}_{0}$ captures the global structure of the signal and lies near the high-density region of $p(\mathbf{x}_{0})$. Stage 2 requires only that $\hat{\mathbf{x}}_{0}$ is close to the data manifold; the specific initialization path is irrelevant, ensuring complete decoupling between the two stages.
Stage 2: Likelihood-Driven Refinement. Given the initialization $\hat{\mathbf{x}}_{0}$, we perform MCMC sampling driven solely by the measurement likelihood. Under the Gaussian measurement model, the likelihood gradient is globally Lipschitz and the log-density is strongly concave, giving the unadjusted Langevin algorithm (ULA) well-established convergence guarantees that depend on the specific type of regularization [10.1111/rssb.12183, AAP1238]. Following ULA, the update rule is
@@
where $\eta$ is the step size and $\boldsymbol{\epsilon}_{j}\sim\mathcal{N}(0,\mathbf{I})$. In standard MCMC, the prior gradient $\nabla_{\mathbf{x}_{0}}\log p(\mathbf{x}_{0})$ serves as a regularization term. However, because the diffusion initialization places $\hat{\mathbf{x}}_{0}$ in a region where $\|\nabla\log p(\mathbf{x}_{0})\|=O(\varepsilon)$, this contribution is numerically negligible. Once we abandon the time-marginal requirement entirely, convergence is governed purely by the geometry of the likelihood, and far fewer steps suffice. We therefore omit the prior gradient and obtain a likelihood-only refinement:
@@
$\mathbf{x}_{0}^{(j+1)}=\mathbf{x}_{0}^{(j)}-\frac{\eta}{\gamma^{2}}\,\nabla_{\mathbf{x}_{0}^{(j)}}\left\|\,\mathbf{y}-\mathcal{A}\!\left(\mathbf{x}_{0}^{(j)}\right)\right\|^{2}+\sqrt{2\eta}\,\boldsymbol{\epsilon}_{j}.$ (9)
This refinement drives the estimate toward the measurement-consistent solution but reduces variability in noise-corrupted or unobserved components. To restore this variability, each refinement cycle is followed by a re-noising step. This controlled stochasticity prevents overfitting and allows the next cycle to explore from a fresh state while maintaining measurement consistency. Unlike maximum a posteriori (MAP) estimation, our method performs stochastic sampling in the diffusion-defined latent space, exploring the posterior rather than collapsing to one mode. To improve convergence stability, we adopt an EDM-inspired annealing schedule in which the step size decays polynomially (exponent in [4,7][-4,-7]), concentrating more iterations in the low-noise regime where fine-scale refinement is most effective.
Algorithm 1 DAPS++ Inference with Decoupled Initialization-Refinement
Require: Measurement $\mathbf{y}$, operator $\mathcal{A}$, effective noise variance $\gamma^{2}$; diffusion score $s_{\theta}$, noise schedule $\{\sigma_{t}\}$, threshold $\bar{\sigma}$, refinement steps $J$, step sizes $\{\eta_{j}\}$, total cycles $K$.
Ensure: Final reconstruction $\hat{\mathbf{x}}_{0}$.
1: Initialize $x_{\mathrm{in}}^{(0)}\sim\mathcal{N}(\mathbf{0},\sigma_{\max}^{2}\mathbf{I})$
2: for $k=1$ to $K$ do $\triangleright$ Stage 1: Diffusion initialization
3:   if $\sigma_{\mathrm{in}}>\bar{\sigma}$ then
4:     $\hat{\mathbf{x}}_{0}^{(k)}\leftarrow x_{\mathrm{in}}^{(k-1)}+\sigma_{\mathrm{in}}^{\,2}s_{\theta}\!\big(x_{\mathrm{in}}^{(k-1)},\sigma_{\mathrm{in}}\big)$
5:   else
6:     $\hat{\mathbf{x}}_{0}^{(k)}\leftarrow\mathtt{ODESolver}\big(x_{\mathrm{in}}^{(k-1)},\sigma_{\mathrm{in}}\big)$
7:   end if $\triangleright$ Stage 2: Likelihood-driven refinement
8:   $z^{(0)}\leftarrow\hat{\mathbf{x}}_{0}^{(k)}$
9:   for $j=1$ to $J$ do
10:     $\mathbf{r}^{(j-1)}\leftarrow\mathbf{y}-\mathcal{A}(z^{(j-1)})$
11:     $\nabla_{z}\mathcal{L}\leftarrow\frac{1}{\gamma^{2}}\mathtt{VJP}_{\mathcal{A}}\big(z^{(j-1)},\mathbf{r}^{(j-1)}\big)$
12:     Sample $\boldsymbol{\xi}_{j}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
13:     $z^{(j)}\leftarrow z^{(j-1)}+\eta_{j}\nabla_{z}\mathcal{L}+\sqrt{2\eta_{j}}\,\boldsymbol{\xi}_{j}$ $\triangleright$ ULA refinement
14:   end for
15:   $\tilde{\mathbf{x}}_{0}^{(k)}\leftarrow z^{(J)}$ $\triangleright$ Re-noising for next iteration
16:   Sample $\boldsymbol{\epsilon}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
17:   $x_{\mathrm{in}}^{(k)}\leftarrow\tilde{\mathbf{x}}_{0}^{(k)}+\sigma_{k}\,\boldsymbol{\epsilon}_{k}$
18: end for
19: return $\hat{\mathbf{x}}_{0}\leftarrow\tilde{\mathbf{x}}_{0}^{(K)}$
3.3 Initialization Quality and Convergence Analysis
To justify this decoupled strategy, we analyze the initialization quality and refinement convergence in turn, showing that once the estimate is near the data manifold, each stage needs to be performed only once.
Initialization quality. At a given noise level $\sigma_{t}$, the Tweedie estimate has already recovered the dominant structure of the signal from the score function: large-scale geometry, color distribution, and semantic content are established. The measurement residual at this point is of order $O(|\mathcal{A}|\sigma_{t})+O(\gamma\sqrt{d_{y}})$, where $d_{y}$ is the measurement dimension. At early annealing stages where $\sigma_{t}$ is large, the residual is also large, but high precision is not required since re-noising at the end of each cycle erases fine-scale corrections; at late stages where $\sigma_{t}$ is small, the residual has shrunk to near the noise floor and only a few Langevin steps are needed. With $\bar{\sigma}=10\gamma$, the transition from Tweedie to a single RK4 step occurs when the first-order truncation error begins to dominate the residual.
Convergence across problem types. The number of refinement steps required in Stage 2 is governed by two properties of the measurement operator in the neighborhood of $\hat{\mathbf{x}}_{0}$. The first is the condition number of the forward map, which controls the convergence rate of the likelihood gradient. The second is the degree of convexity of the negative log-likelihood near the initialization, which determines whether local refinement can reach the correct solution. A formal convergence bound relating these quantities to the required step count is provided in the Appendix. For well-conditioned linear operators, the likelihood is quadratic and convergence is fast: super-resolution requires only 2 steps because the downsampling operator preserves most signal energy, while Gaussian deblurring requires more steps due to its broader eigenvalue spread. For mild nonlinearities such as per-pixel tone mapping in HDR, the likelihood remains approximately convex near the initialization and convergence is comparable to the linear case (5 steps). For stronger nonlinearities such as nonlinear deblurring, the signal-dependent kernel introduces inhomogeneous curvature, requiring more iterations (50 steps). In phase retrieval, the likelihood has multiple equivalent minima; if the initialization falls in the wrong basin, local refinement cannot recover the correct solution.